Today jack is twice as old as Jill In ten years time jack will be one and a half times as old as Jill ,how old will jack and Jill be in twenty years?
Answer by stanbon(75887)

Today jack is twice as old as Jill In ten years time jack will be one and a half times as old as Jill ,how old will jack and Jill be in twenty years?
-----------------------------
J = 2L
J+10 = (3/2)(L+10)
------
Substitute for "J" and solve for "L":
2L+10 = (3/2)L + 15
-----
(1/2)L = 5
Jill = 10 yrs
----
Jack = 2L = 20 yrs
-----------------
Cheers,
Stan H.
------------